2. Solve (i) Find the general solution of y + y = 0. Here, dont have to check the Wronskain not zero for the fundamental s
Answer #1

we have Page 1 @ y #y = 0 - 0 or 3 you becomes - $340) = 0 ( 3 + y =o 2²€ AI) = 0 D2=0 or P + ZO D=0,0,– 1 then solution ofI we have y Initial condition = 0 - i Je = sine ye - e - ri -01) Now from Yy = @ enged - @ yu = C,+Garces 1 Now Applying thehey dear! This is your complete solution in detail. Plz like it ???I did hard work for you.